ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 16, 20 Marcus, Andrei, “Babeş-Bolyai” University Cluj, Romania Derived equivalences and Dade’s conjectures We show that certain derived equivalences between blocks of finite group algebras preserve the invariants involved in the various forms of Dade’s conjecture. It will follow that in the abelian defect group case, an extended form of Broué’s conjecture implies Dades’s inductive conjecture. References: 1. E. C. Dade, Counting characters in blocks 2.9, Proc. of the 1995 Conference On Representation Theory at Ohio State University.

2. A. Marcus, On equivalences between blocks of group algebras: reduction to the simple components, J. Algebra 184 (1996), 372–396.